OpenStudy (anonymous):

The area of a football field is x2 + x - 42. Find the length and width of the field in terms of x.

OpenStudy (anonymous):

i think just solve for x an the larger value for x is the length and the smaller one is the width. you need to find two numbers that when mupltied together = -42 and when added together = 1. (factoring)

OpenStudy (anonymous):

So It Would Be (x - 7) and (x + 6)

OpenStudy (anonymous):

yeah but now that i look at that you can't have a negative length or width. ummm. forget what i said. give me a second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay um. A = L*W = x^2+x-42 so the only thing i can think of to write the length in terms of x and the width in terms of x is to go L=( x^2+x-42)/W then do the same for the width. I'm not sure if that's right but that's the best i got

OpenStudy (anonymous):

ok thanks tho